# Tensor networks for quantum computing

### v1

@@ Acknowledgments
A. B., R. E., V. K., C. M., A. M., F. N., D. M., M. Pere., M. Pf., and V. V. were supported by Terra Quantum AG. A. N. was supported by Natural Sciences and Engineering Research Council of Canada (NSERC) Alliance Quantum Program (Grant ALLRP-578555), and the Canada First Research Excellence Fund, Quantum Materials and Future Technologies Program. S.Ṁ. is partially supported by the Prime Contract No. 80ARC020D0010 with the NASA Ames Research Center and acknowledges funding from DARPA under IAA 8839. A. B. thanks Ilia Luchnikov, Mikhail Litvinov and Stefanos Kourtis for valuable discussions. A.A., Z.H., M.L., M.A.P., R.S. and M.P. thank their colleagues at the Global Technology Applied Research center of JPMorganChase for support and helpful discussions.

### v2

@@ Acknowledgments
A. B., R. E., V. K., C. M., Al. M., Ar. M., V. M., D. M., F. N., M. Pere., M. Pf. and V. V. were supported by Terra Quantum AG. A. N. was supported by Natural Sciences and Engineering Research Council of Canada (NSERC) Alliance Quantum Program (Grant ALLRP-578555), and the Canada First Research Excellence Fund, Quantum Materials and Future Technologies Program. S.Ṁ. is partially supported by the Prime Contract No. 80ARC020D0010 with the NASA Ames Research Center and acknowledges funding from DARPA under IAA 8839. A. B. thanks Ilia Luchnikov, Mikhail Litvinov and Stefanos Kourtis for valuable discussions. A. A., Z. H., A. K., M. L., M. Perl., M. Pi., M. S. and R. S. thank their colleagues at the Global Technology Applied Research center of JPMorganChase for support and helpful discussions.
Author contributions
Yu. A., A. B., M. L. and V. V. conceptualized the work. A. B., R. E., J. G., R. H., A. K., M. L., A. N., F. N. and V. V. contributed to the Tensor-network methods section. M. L., A. N., F. N. contributed to the Simulation of quantum computation section. A. A., R. H., Z. H., A. K., M. L., M. Pere., V. V. contributed to the Quantum circuit synthesis section. A. B., R. H., M. Perl., M. S., B. V. contributed to the Quantum error correction and mitigation section. A. A., V. K. contributed to the Tensor networks for quantum machine learning section. All authors contributed to the Introduction and the Discussion and outlook sections. All authors were involved in shaping the direction of the manuscript, as well as in its discussion, review, and editing.